Simplify each expression by rewriting it as a single logarithm...


2 log a-3 log b=


x ln m+ y ln n=
1
Expert's answer
2011-05-17T08:12:13-0400
2 log a - 3 log b = log a2/b3
x ln m + y ln n = ln (xm yn)